OpenStudy (legit_sarahh):

whats the solution to the system of equations. x+y=2 y=2x-1

OpenStudy (dtan5457):

X=1 y=1 How do you find this? They already tell you that y=2x-1 so plug that into x+y=1 which is x+2x-1=2 Simplify that and you get 3x=3 x=1, plug that back to 1+y=2 y=1

OpenStudy (legit_sarahh):

thanks